For all of the problems below, here is the reaction: A <=> B. Assume in each case that the temperature is 300 K, except as noted.

1. What is the Delta G0' for a reaction if the Delta G is -20 kJ/mol when there is 5 times as much A as B?

2. When the Delta G for a reaction is -10 kJ/mol and the Delta G0' is -10 kJ/mol, what is the ratio of A to B?

3. If the ratio of A/B in #1 stays the same, what happens to the Delta G when the temperature drops 50 degrees?

4. A reaction is favorable when there is twice as much A as B and also when there is twice as much B as A. Is this possible? If not, why not. If so, explain.
Answer #1

1)

Q = B / A = 1 / 5 = 0.2

ΔG = ΔG∘ + RT ln Q

- 20 = ΔG∘ + 8.314 x 10^-3 x 300 x ln 0.2

ΔG∘ = - 15.98 kJ/mol

2)

ΔG = ΔG∘ + RT ln Q

-10 = - 10 + RT ln [B]/[A]

B / A = 1

A / B = 1

3)

ΔG = ΔG∘ + RT ln Q

if A/B = 1

Q = B/A = 1

then lnQ = 0

so

ΔG = ΔG∘

there is no change when temperature drops to 50 C

4)

ΔG = ΔG∘ + RT ln Q

yes favorable